OpenStudy (anonymous):

2y-5x=6 in graphically form

OpenStudy (anonymous):

if x =0, y =?

OpenStudy (anonymous):

if y =0, x =?

OpenStudy (thesecret20111):

Set x to 0 and solve for y. Then set y to 0 and solve for x

OpenStudy (mathmale):

It'd be helpful to all of us, including you (shinnG) if you'd specify what your goal is here. For example: GRAPH 2y-5x=6 (your "2y-5x=6 in graphically form" is not properly presented in terms of English). Try again. Starting with 2y-5x=6, ... what do you wish to do ... ?